OpenStudy (anonymous):

Let f(x) = x - 2 and g(x) = x2 - 7x - 9. Find f(g(-1)).

OpenStudy (anonymous):

Solve for g(-1) first what do you get?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

now plug that into f(x)

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

check your answer for g(x)

OpenStudy (anonymous):

sorry but i keep making the same mistake

OpenStudy (anonymous):

That's how we learn. I do the same thing

OpenStudy (anonymous):

well i got -3 again

OpenStudy (anonymous):

(-1) squared = 1 -7 * -1 = 7 -9 = -9 Thus g(x) = -1 Correct?

OpenStudy (anonymous):

Okay i see ,i was messing up on squaring -1

OpenStudy (anonymous):

Good job..When working functions embedded in other functions solve from the inside to the outside.